Have there been any problems solved such as .. Solving a linear equation with several occurrences of the variable Problem type 4?
-2(4x -8)+4x = 6(x + 3)

I stuggle with the changing of the signs and this causes me to get the wrong answer every time. Can someone please explain when to change the sign in these kind of problems. If any of the Tutors can help it would be grately appreciated
I have 5 weeks of Algebra and I don't want to fail the class again. I'm in week one now
Thanking you in advance.

Answer by mananth(16946)

-2(4x -8)+4x = 6(x + 3)
...
Please remember in short negative * negative = positive ( -2*-2 =+4)
negative * positive = negative (-2*+2=-4)
-8x+16 +4x=6x+18
whenever you are having terms added & subtracted in an equation like above
your aim is to bring the terms with x on one side and the numbers to the other side, to find the value of x.
-8x+4x-6x= -16+18. ( note that 6x & +16 have changed sides in the equation hence their signs have beeeen changed.
-10x=+2
divide by -10
-10x/-10 = 2/-10
x= 2/-10
x=-1/5
....